OpenStudy (anonymous):

expand: log 3square root 0f x

OpenStudy (p0sitr0n):

log3sqrt(x) ?

OpenStudy (lgbasallote):

\(\Huge \log_{3} x^{\frac{1}{2}}\) that give you any ideas??

OpenStudy (p0sitr0n):

1/2log3x

OpenStudy (anonymous):

no log and the symbol root with 3

OpenStudy (p0sitr0n):

cubic root?

OpenStudy (p0sitr0n):



OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah

OpenStudy (p0sitr0n):

ok so logx^1/3 move 1/3 as coefficient gives 1/3(logx)

OpenStudy (anonymous):

where did you get the 1/3

OpenStudy (anonymous):

because \(\sqrt[3]{x}=x^{\frac{1}{3}}\) and \(\log(x^r)=r\log(x)\)

OpenStudy (anonymous):

ok that helps soo much!! thnxs

OpenStudy (anonymous):

yw